OpenStudy (baby456):

Fan +Medal help with factoring x^2-x-72

OpenStudy (misssunshinexxoxo):

x^2-x-72 =x^2-9x+8x-72 x(x-9)+8(x-9) (x-9)(x+8)

OpenStudy (baby456):

still cconfused

OpenStudy (misssunshinexxoxo):

Ok do you know how to factor?

OpenStudy (anonymous):

(x+8)(x-9)

OpenStudy (misssunshinexxoxo):

@baby456 do you need more explanation?

OpenStudy (baby456):

yes

OpenStudy (anonymous):

look at the middle section realise that -x is the same as -9x+8x do you get that part? @baby456

OpenStudy (baby456):

no

OpenStudy (anonymous):

then try thinking of it without the x -1 = -9+8 now do you understand? @baby456